[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon Aug 14, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Aug 14 20:06:31 UTC 2017


HSU Shift Log for Mon Aug 14 2017 16:06:25 GMT-0400 (EDT)

Shift: H

Forecaster: D Brown

Technical Issues Summary:

ATCF issues: None

N-AWIPS issues: None

AWIPS-2/NCP issues: None

Other Techincal issues: None

Miscellaneous issues: An FYI- if the background on the NAWIPS workstations  
changes colors, try ctrl-alt-backspace to restart the X server (resets the  
graphics) This will log you out.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170814/dbe6899f/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list